OpenStudy (anonymous):

64^4x=16^x+4 Please Help!

OpenStudy (anonymous):

\[64^4x=16^{x+4}\] is that it?

OpenStudy (anonymous):

no i mean\[64^{4x}=16^{x+4}\]

OpenStudy (anonymous):

divide both sides by 16^x+4: \[4^{10x-8}=1\]

OpenStudy (anonymous):

take log base 4 of both sides so you get: \[10x-8=0\] and now you can solve it by yourself

OpenStudy (anonymous):

you can also solve it by writing it as : \[4^{3*4x}=4^{2(x+4)}\] then: \[3*4x=2(x+4)\] and solved as normal

OpenStudy (anonymous):

Thanks a bunch I was stuck on that one for awhile